# LLMs for Generalizable Language-Conditioned Policy Learning under Minimal Data Requirements

### v1

@@ 5.4 Q4: Data efficiency and Scaling of TEDUO
Table 4: Performance vs. compute power / number of training goals.
@@
Impact of state abstraction. In Figure 4, we look at the performance of the learned Q-learning policies (i.e. the policies $\{\pi_{g}\}_{g\in{\mathcal{G}}^{tr}}$ obtained at the end of TEDUO step 1+2 for different sizes of observational dataset $\mathcal{D}$. This experiment is realized with and without the ablation of the abstraction function during step 1. As anticipated, the efficacy of the learned policies improves with the increasing size of the dataset $\mathcal{D}$ until reaching a plateau. On average, our LLM-based state abstraction function reduces the number of unique states within each abstract MDP by  10% (Fig. C.8 in the Appendix). Due to the reduction in state space size, the abstraction function significantly enhances the data efficiency of our training method across all three performance metrics. Furthermore, the size of the state spaces $\mathcal{S}_{\phi}^{g}$, corresponding to the subset of features relevant for identifying the completion of the goal is reduced to just around 20% of the original state space size (Fig. C.8 in the Appendix). This reduces the size of the goal detection datasets for training and the subsequent goal-identification described in section 3.1.2 5-fold.
Compute power is the new bottleneck. Given a fixed observational dataset ${\mathcal{D}}$, we can expand at no extra cost the fine-tuning dataset $\mathcal{D}^{SFT}$ by introducing more training goals in ${\mathcal{G}}^{tr}$. Yet, larger $\mathcal{D}^{SFT}$ necessitates more compute power for training the LLM agent. Table 5.4 demonstrates the scaling of our method with compute power. As expected, training on a wider range of goals results in an improved performance on unseen test goals. We do not observe a plateau in performance metrics, suggesting that with additional compute further gains may be possible. Consequently, our approach shifts the bottleneck from the limited availability of real observational data to computational power.
@@ 6 Discussion
Limitations. Leveraging LLMs’ prior knowledge enables efficient policy generation with minimal data. However, some applications may benefit more than others. First, certain scenarios may be out of distribution even for LLMs trained on extensive Internet data. Second, we assume that the environment state can be represented textually, which, although feasible for many applications due to language’s expressiveness, may not be ideal in all cases. Third, due to the discrete nature of LLM tokenization, using fine-tuned LLMs to directly output actions requires discretization of the action space, which can hinder performance in continuous control tasks. Lastly, while data requirements are minimal, they still assume some practitioner knowledge of the environment and the data ${\mathcal{D}}$ to propose training goals ${\mathcal{G}}^{tr}$ likely achievable in ${\mathcal{D}}$ (see Appendix C.1 for details).
Conclusions. We introduced a novel framework for training natural language instruction-following agents capable of generalizing to new states and instructions in a zero-shot setting. This is the first method to learn natural language goal-conditioned policies in an offline setting using unlabeled data and it surpasses the current state-of-the-art in zero-shot generalization for both goal and domain adaptation. To achieve this, TEDUO teaches environment dynamics to an LLM, enabling it to interact effectively within the given environment—a task where out-of-the-box LLMs typically underperform. Given the demonstrated flexibility of LLMs in generalizing to new goals and environments, future work could explore the potential of this approach in learning across multiple environments with distinct action and state spaces.
Reproducibility statement. The environment used to analyze and benchmark the methods are publicly available. Within the TEDUO pipeline and for benchmarking we use open-source language models. Every step of our method is clearly stated in Section 3. Additional details to reproduce the experiments, including prompts and implementation of baselines, are presented in Appendix B. The code for this paper is provided as supplementary material.
@@ A.1 Generalization in offline Reinforcement Learning
Language-conditioned RL. Our work addresses the problem of goal-conditioned RL, where goals are expressed in natural language. While using language to specify goals is natural and broadens the range of possible goals it comes with the challenge of grounding the semantics of language in the environment state space and dynamics. Such language-instruction following agents have been widely studied in both reinforcement learning and imitation learning contexts. However, most existing methods either rely on access to an online environment for interaction (Fu et al., 2018; Bahdanau et al., 2018; Jiang et al., 2019; Mirchandani et al., 2021) or require costly, goal-annotated expert datasets of offline demonstrations (Stepputtis et al., 2020; Lynch & Sermanet, 2021; Xiao et al., 2023; Brohan et al., 2023b; a). In contrast, our approach does not assume any environment-provided reward signal or access to real-time exploration. Furthermore, in terms of generalization to new natural language instructions, we distinguish between evaluation on instructions that simply paraphrase the training goals (Nair et al., 2022; Lynch & Sermanet, 2021) from those that represent semantically novel goals. Similar to the works of Xiao et al. (2023); Brohan et al. (2023a); Stepputtis et al. (2020); Shridhar et al. (2021a); Jang et al. (2022), our focus is on the latter, more challenging scenario.
Domain Generalization. While the previous section addressed generalization to new goals, this section focuses on generalization to novel state-action transitions. This type of generalization extends beyond goal-conditioned RL, as it is essential even for single-goal RL. It has been widely studied and observed that Offline RL methods often overfit to the training distribution of state-action transitions, resulting in poor performance when the test distribution differs. Various approaches have been proposed to address this distribution shift, including regularization techniques Kostrikov et al. (2021); Kumar et al. (2020), model-based RL Yu et al. (2020); Kidambi et al. (2021), and enhanced representation learning Mazoure et al. (2021); Fan & Li (2022). In TEDUO, we intentionally avoid domain generalization when solving the abstract MDPs in step 2 to prevent providing incorrect examples to the LLM in step 3. However, our method achieves domain generalization by leveraging the zero-shot capabilities of the fine-tuned LLM. Future work could enhance TEDUO by replacing tabular Q-learning in step 2 with a method that generalizes to new state-action transitions.
@@
Hindsight labeling. Labeling data for goal-conditioned RL requires the design of reward functions for each goal. The most common approach for designing the rewards relies on handcrafted methods that are often require multiple refinements through trial and error (Knox et al., 2022). With a large number of goals, manual reward design becomes infeasible. Inverse Reinforcement Learning (Ziebart et al., 2008; Fu et al., 2018) attempts to generate reward functions directly from data, but it requires a large amount of expert demonstrations. Recent studies have explored the use of LLMs and VLMs as reward functions. These methods typically involve creating a preference dataset (Klissarov et al., 2023), comparing the cosine similarity between natural language goals and state representations, or leveraging the coding abilities of LLMs (Yu et al., 2023a), especially in an online iterative fashion (Ma et al., 2023; Xie et al., 2024). These approaches are however aimed at densifying the reward signal. In contrast, our method requires generating reward labels for a large number of goals (approx. 100-1000), making the scalability of the process crucial. Therefore, we focus on generating rewards with a limited number of LLM calls. Our approach relies on LLM-based detection of task completion, which has been proven effective by Kwon et al. (2023). We further reduce the number of LLM calls by approximating the LLM-generated rewards with lightweight proxy neural networks.
@@ A.3 Large Language Models for Decision Making
General-purpose LLMs for decision making. Utilizing off-the-shelf LLMs has gained significant attention due to its simplicity. In decision-making, LLMs have been used to create assistance functions within training pipelines to enrich data (Klissarov et al., 2023; Yu et al., 2023a; Ma et al., 2023; Xie et al., 2023; Laskin et al., 2022), and as high-level planners during inference to guide traditional RL policies (Shah et al., 2023; Ahn et al., 2022). Additionally, there has been growing interest in using general-purpose LLMs directly as decision-making agents (Yao et al., 2023). Improving the reasoning capabilities of LLM agents is now an active research area, focusing on methods that are independent of traditional RL. These include iterative prompting techniques such as self-reflection (Ji et al., 2023), chain of thought reasoning (Wei et al., 2023), and integration with planning algorithms like Monte Carlo Tree Search (Pouplin et al., 2024). Nevertheless, such methods have been shown inefficient in completing complex, multi-step decision-making tasks in dynamic environments (Finn, 2024; Szot et al., 2024). To effectively use the knowledge embedded in LLMs for solving RL problems, these models need to be grounded in the dynamics of the environment.
@@
A.4 Testing Environments

### v2

@@ 5.4 Q4: Data efficiency and Scaling of TEDUO
Impact of state abstraction. Figure 5 shows the performance of the Q-learning policies (i.e., policies $\pi_{g}$ obtained at the end of TEDUO step 1+2) against the size of the observational dataset $\mathcal{D}$. This experiment is realized with and without the abstraction function during step 1. As anticipated, the efficacy of the learned policies improves with increasing size of the dataset $\mathcal{D}$. On average, the LLM-based state abstraction reduces the number of unique states by 10% (see Fig. B.2). The reduction in state space size significantly enhances data efficiency of our training method across all three performance metrics. Furthermore, the size of the state spaces $\mathcal{S}_{\phi}^{g}$, corresponding to the subset of features relevant for identifying the completion of the goal is reduced to just around 20% of the original state space size (Fig. B.2 in the Appendix). This reduces the size of the datasets used for training and the subsequent goal-identification 5-fold.
Compute power is the new bottleneck. Given a fixed observational dataset ${\mathcal{D}}$, we can expand at no extra cost the fine-tuning dataset $\mathcal{D}^{SFT}$ by introducing more training goals in ${\mathcal{G}}^{tr}$. Yet, larger $\mathcal{D}^{SFT}$ necessitates more compute power for training the LLM agent.
Table 4: Performance vs. compute power.
@@
Table 4 demonstrates the scaling of our method with compute power. As expected, training on a wider range of goals results in an improved performance on unseen test goals. We do not observe a plateau in performance metrics, suggesting that with additional compute further gains may be possible. Consequently, our approach shifts the bottleneck from the limited availability of real observational data to computational power.
@@ 6 Discussion
Limitations. Leveraging LLMs’ prior knowledge enables efficient policy generation with minimal data. However, some applications may benefit more than others. First, certain scenarios may be out of distribution even for LLMs trained on extensive Internet data. Second, we assume that the environment state can be represented textually, which, although feasible for many applications due to language’s expressiveness, may not be ideal in all cases (this can be owever mitigated by employing VLMs which we leave as future work). Third, due to the discrete nature of LLM tokenization, using fine-tuned LLMs to directly output actions requires discretization of the action space, which can hinder performance in continuous control tasks. Lastly, while data requirements are minimal, they still assume some practitioner knowledge of the environment and the data ${\mathcal{D}}$ to propose the list of goals ${\mathcal{G}}$ (see Appendix B.2 for details).
Conclusions. We introduced a novel framework for training natural language instruction-following agents capable of generalizing to new states and instructions in a zero-shot setting. This is the first method to learn natural language goal-conditioned policies in an offline setting using observational data that is both unlabeled and non-expert. The success of TEDUO relies on the synergy between the mutual strengths of conventional RL and LLMs. Given the demonstrated flexibility of this method in generalizing to new goals and environments, future work could explore its potential in learning across multiple environments with distinct action and state spaces.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab. Thomas Pouplin’s research is supported by funding from AstraZeneca. Katarzyna Kobalczyk’s research is supported by funding from Eedi. Hao Sun’s research is supported by funding from the Office of Naval Research (ONR).
@@ A.1 Generalization in offline Reinforcement Learning
Language-conditioned RL. Our work addresses the problem of goal-conditioned RL, where goals are expressed in natural language. While using language to specify goals is natural and broadens the range of possible goals it comes with the challenge of grounding the semantics of language in the environment state space and dynamics. Such language-instruction following agents have been widely studied in both reinforcement learning and imitation learning contexts. However, most existing methods either rely on access to an online environment for interaction (Fu et al., 2018; Bahdanau et al., 2018; Jiang et al., 2019; Mirchandani et al., 2021) or require costly, goal-annotated expert datasets of offline demonstrations (Stepputtis et al., 2020; Brohan et al., 2023b, a). In contrast, our approach does not assume any environment-provided reward signal or access to real-time exploration. Furthermore, in terms of generalization to new natural language instructions, we distinguish between evaluation on instructions that simply paraphrase the training goals (Nair et al., 2022; Lynch & Sermanet, 2021) from those that represent semantically novel goals. Similar to the works of Xiao et al. (2023); Brohan et al. (2023a); Stepputtis et al. (2020); Shridhar et al. (2021a); Jang et al. (2022), our focus is on the latter, more challenging scenario.
Domain Generalization. While the previous section addressed generalization to new goals, this section focuses on generalization to novel state-action transitions. This type of generalization extends beyond goal-conditioned RL, as it is essential even for single-goal RL. It has been widely studied and observed that Offline RL methods often overfit to the training distribution of state-action transitions, resulting in poor performance when the test distribution differs. Various approaches have been proposed to address this distribution shift, including regularization techniques (Kostrikov et al., 2021; Kumar et al., 2020), model-based RL (Yu et al., 2020; Kidambi et al., 2021), and enhanced representation learning (Mazoure et al., 2021; Fan & Li, 2022). In TEDUO, we intentionally avoid domain generalization when solving the abstract MDPs in step 2 to prevent providing incorrect examples to the LLM in step 3. However, our method achieves domain generalization by leveraging the zero-shot capabilities of the fine-tuned LLM. Future work could enhance TEDUO by replacing tabular Q-learning in step 2 with a method that generalizes to new state-action transitions.
@@
Hindsight labeling. Labeling data for goal-conditioned RL requires the design of reward functions for each goal. The most common approach for designing the rewards relies on handcrafted methods that are often require multiple refinements through trial and error (Knox et al., 2022). With a large number of goals, manual reward design becomes infeasible. Inverse Reinforcement Learning (Ziebart et al., 2008; Fu et al., 2018) attempts to generate reward functions directly from data, but it requires a large amount of expert demonstrations. Recent studies have explored the use of LLMs and VLMs as reward functions. These methods typically involve creating a preference dataset (Klissarov et al., 2023), comparing the cosine similarity between natural language goals and state representations, or leveraging the coding abilities of LLMs (Yu et al., 2023a), especially in an online iterative fashion (Ma et al., 2023; Xie et al., 2024). These approaches are however aimed at densifying the reward signal. In contrast, our method requires generating reward labels for a large number of goals (approx. 100-1000), making the scalability of the process crucial. Therefore, we focus on generating rewards with a limited number of LLM calls. Our approach relies on LLM-based detection of task completion, which has been proven effective by (Kwon et al., 2023). We further reduce the number of LLM calls by approximating the LLM-generated rewards with lightweight proxy neural networks.
@@ A.3 Large Language Models for Decision Making
General-purpose LLMs for decision making. Utilizing off-the-shelf LLMs has gained significant attention due to its simplicity. In decision-making, LLMs have been used to create assistance functions within training pipelines to enrich data (Klissarov et al., 2023; Yu et al., 2023a; Ma et al., 2023; Xie et al., 2023; Laskin et al., 2022), and as high-level planners during inference to guide traditional RL policies (Shah et al., 2023; Ahn et al., 2022). Additionally, there has been growing interest in using general-purpose LLMs directly as decision-making agents (Yao et al., 2023b). Improving the reasoning capabilities of LLM agents is now an active research area, focusing on methods that are independent of traditional RL. These include iterative prompting techniques such as self-reflection (Ji et al., 2023), CoT reasoning (Wei et al., 2023), and integration with planning algorithms like Monte Carlo Tree Search (Pouplin et al., 2024). Nevertheless, such methods have been shown inefficient in completing complex, multi-step decision-making tasks in dynamic environments (Finn, 2024; Szot et al., 2024). To effectively use the knowledge embedded in LLMs for solving RL problems, these models need to be grounded in the dynamics of the environment.
@@
Appendix B Additional Results
B.1 Webshop results
Main Results. The results averaged over 200 instructions are shown in Table B.1. Unlike Minigrid, Webshop does not feature multiple environment types, prohibiting generalization evaluation across environments. The scores represent the true environment reward (scaled to [0, 100]).
Table B.1: Online evaluation of generalization performance for the Webshop environment. Results averaged over 200 instructions.
{tblr}
width = colspec = Q[410]Q[187]Q[135]Q[200], cells = c, hline1-2,4,6 = -, Method & Goals Score Episode Length
ReAct-Llama-3-8B Training/Testing 8.4 (±0.8) 14.1 (±0.1)
ReAct-Llama-3-70B Training/Testing 13.8 (±0.9) 14.2 (±0.1)
TEDUO-Llama-3-8B Training 52.1 (±0.6) 6.6 (±0.1)
TEDUO-Llama-3-8B Testing 44.4 (±0.6) 6.9 (±0.1)
This benchmark demonstrates that TEDUO successfully learns from unlabeled trajectories in the Webshop environment. TEDUO significantly outperforms the ReAct prompting approach, particularly when applied to smaller models (compared to GPT-3.5 in the original paper). While ReAct prompting provides some improvement over standard prompting, its performance remains limited, whereas TEDUO achieves substantially better results.
Table B.2: Ablation study for the Webshop environment. Results averaged over 200 instructions.
{tblr}
width = colspec = Q[410]Q[187]Q[135]Q[200], cells = c, hline1-2,4,6 = -, Method & Goals Score Episode Length
Data collection policy (LLM-random) Training/Testing 5.6 (±0.5) 12.3 (±0.2)
Step 1 2 (GCRL) Training 56.2 (±0.6) 6.9 (±0.1)
TEDUO-Llama-3-8B Training 52.1 (±0.6) 6.6 (±0.1)
TEDUO-Llama-3-8B Testing 44.4 (±0.6) 6.9 (±0.1)
Ablation study: Similar to the approach in the paper, table B.2 provides an ablation study below that summarizes the performance improvements achieved after each TEDUO step. The results demonstrate that TEDUO Steps 1 & 2 effectively produce improved policies over the data collection policy for the training goals. Furthermore, TEDUO Step 3 successfully achieves its generalization objective, extending the learned policy’s performance to new instructions. Compared to the BabyAI environment, the fine-tuned LLM does not outperform the Q-learning policies trained on the training instructions. This limitation can be attributed to the lack of diversity in the initial states. In the BabyAI environment, agents can start from any grid position, including configurations unknown to the tabular Q-learning policies, often resulting in failure trajectories. In contrast, the fine-tuned LLM demonstrates successful generalization in such cases. However, the webshop environment presents a single initial state—the search bar webpage. To surpass the Q-learning policies in this setting, the LLM would need to contradict its learned behaviors, making superior performance unattainable.
B.2 The impact of the data collection policy
To empirically analyze this, we consider two randomized policies:
  • B) Random policy: Actions are drawn uniformly at random from the action space. This policy represents an agent that explores the environment without a specific goal. Although this scenario is less common in real-world settings—where agents typically pursue objectives—it remains applicable to batch RL, particularly when learning from untrained agents with no prior knowledge.
Figure B.1: Impact of data collection policy. The x-axis shows the proportion of data ${\mathcal{D}}$ collected with policy A vs. policy B for a fixed size of ${\mathcal{D}}$. a) The y-axis shows $|\mathcal{S}_{0}^{g}|$, i.e. the number of unique initial abstract states $s_{0}^{g}$ for which $g$ is reachable with the learned policy $\pi^{g}$. Values are min-max normalized across all 5 mixture policies. b) The y-axis shows the same values as plot a), averaged across 14 goals, bars represent the standard error.
The metric of fidelity. The likelihood of the final fine-tuned LLM in solving a given goal $g$ from any starting position is directly linked to the fidelity of the corresponding Q-learning policy, $\pi^{g}$. Thus, we chose to measure the fidelity of our starting point inputs by computing the number of initial abstract states $s_{0}^{g}$, from which $g$ can be solved using $\pi^{g}$. We define a goal $g$ as reachable from $s_{0}^{g}$ if by iteratively selecting the actions $a_{t}^{*}=\arg\max\pi^{g}(\cdot|s_{t}^{g})$, the agent can eventually reach $g$. We denote this number of states with $|\mathcal{S}^{g}_{0}|$. The set of candidate initial states to compute $|\mathcal{S}_{0}^{g}|$ is the entire abstract state space.
Figure B.1 illustrates that the optimal data collection policy varies by goal. For some goals policy A works better, while for others it is the fully random policy. Importantly, a comparable amount of synthetic action sequences for fine-tunning the LLM can be extracted using either policy A or B. Averaging across all goals, we find that policy A tends to perform better for goals in ${\mathcal{G}}_{\pi}$ than those not in ${\mathcal{G}}_{\pi}$.
Explanation: We note that only a subset of initial states $s_{0}^{g}$ has been used during data collection. Using a goal-oriented policy to collect trajectories results in limited exploration of the state space. Consequently, the learned policies $\pi^{g}$ are expected to excel at solving $g$ when started from a state that has been visited in $\mathcal{D}$, but are not guaranteed to succeed when started from $s_{0}^{g}$ not visited in $\mathcal{D}$. This is why, including more exploration during data collection can help in ensuring that the learned policies can reach their respective goals from anywhere in the state space. However, relying solely on random exploration is only effective for simple goals that are likely to be reached by chance. Instead, in our main experiments, we use a goal-oriented collection policy that uniformly samples across goals of varying difficulty. This results in a more diverse dataset that includes complex behaviors and better reflects realistic settings, where completely random agents don’t exist.
Future work could explore optimizing the set of training goals $\mathcal{G}^{tr}$ to maximize the alignment of $\mathcal{G}^{tr}$ with a given dataset $\mathcal{D}$. Yet, the necessity of aligning ${\mathcal{D}}$ and $\mathcal{G}^{tr}$ is moderated by two factors. First, as shown in subsection 5.4, the abstraction function reduces the complexity of the abstract MDPs, requiring fewer data samples. Second, since extending the list of goals in ${\mathcal{G}}^{tr}$ is computationally inexpensive, we can continually seek better alignment.
B.3 Abstraction function
Figure B.2 presents the performance of the abstraction function used in the BabyAI environment.
Figure B.2: Reduction in count of unique states due to applying the $\mathrm{LLM}(g)$ abstraction functions and the relative size of the reduced abstract feature space ${\mathcal{S}}^{g}_{\phi}$, containing only features necessary to identify the completion of a goal $g$.
B.4 TEDUO step 2: Deep Q-Learning.
The step 2 of the TEDUO framework is independent of the specific offline reinforcement learning algorithm used to solve the abstract MDPs generated in step 1. Each abstract MDP is represented by a labeled transition dataset $\mathcal{D}^{g}$, consisting of tuples $(s^{g},a,s^{\prime g},r^{g})$, making it compatible with any standard offline RL algorithm. In the implementation presented in Section 3, we used tabular Q-learning as it was most suitable for the environment under consideration. However, to illustrate the scalability of the approach, we show in table B.3 the performance of Deep Q-learning (DQN) (Mnih et al., 2013), demonstrating that it can effectively replace tabular Q-learning in our framework. The deep learning models for Q-value estimation combine CNN and dense layers, consistent with the lightweight models used for reward shaping (see Appendix C.5).While tabular Q-learning is limited to discrete state and action spaces and becomes impractical as their dimensions grow, DQN handles continuous spaces and scales effectively without such constraints.
Table B.3: Ablation study with DQN. Results averaged over 400 $(g,s_{0}^{g})$ pairs.
Method Success Rate [%] Episode Length Invalid Actions [%]
Step 1 + GCBC 7 (±0.6) 474 (±2.3) 11 (±0.1)
Steps 1 & 2 (GCRL- Tabular) 16 (±0.8) 430 (±3.9) 10 (±0.1)
Steps 1 & 2 (GCRL- DQN) 15 (±1.2) 446 (±5.9) 15 (±0.1)
All steps Llama-3-8B 65 (±1.4) 203 (±6.7) 21 (±0.7)
B.5 Reward shaping evaluation
Table B.4: Reward Shaping Benchmark. The accuracy, precision, and recall metrics are computed with a classification threshold ensuring at least 95% precision.
{tblr}
Go to a box
0.90 89 96 38
Pick up a ball 0.75 98 95 83
Open a door 0.92 85 95 85
Go to red door 0.98 94 100 0.2
Go to the tile (5,6) 1.0 100 100 100
Put a box next to a blue ball 0.64 100 100 25
Table B.4 shows the performance of $R_{\theta}(\cdot;g)$ for various types of goals compared to ground truth rewards. The benchmark setup is consistent with the main experiments; details are provided in the Appendix C.5. All goals achieve 95% precision, a crucial metric since false positives lead to generating incorrect data points for ${\mathcal{D}}^{SFT}$ in TEDUO’s step 2. Conversely, false negatives only reduce data points in ${\mathcal{D}}^{SFT}$, which is less critical given our synthetic data abundance (see Section 5.4). Performance varies across goals; for instance, ”go to the red door” has low recall (0.2%), likely due to limited positive examples in the dataset. Expanding the dataset could improve such outcomes.
Table B.5: LLM-only Reward Shaping Benchmark. Accuracy, precision, and recall are computed with respect to LLM-generated labels.
{tblr}
width = colspec = Q[394]Q[196]Q[196]Q[150], cells = c, hline1-2,8 = -, Goals & Accuracy (%) Precision (%) Recall (%)
Go to a box
100 100 100
Pick up a ball 100 100 100
Open a door 100 100 100
Go to a green door 99 88 100
Go to the tile (5,6) 100 100 100
Put a box next to a blue ball 100 100 100
Table B.5 reports the performance of LLM-only reward shaping across a variety of goal types. The model achieves near-perfect accuracy, precision, and recall, indicating a strong ability to generate consistent goal labels from environment states. The slight drop in precision for the Go to a green door task highlights occasional LLM hallucinations, though such cases are rare. Consequently, most residual inaccuracies in the overall reward shaping pipeline come from lightweight downstream models used to reduce computational overhead, rather than from the LLM-generated annotations themselves.
B.6 Benchmark results per goal category
Table B.6: Online evaluation of generalization performance split per goal category. This is the success rate [%] presented in Table 1 with the 400 $(g,s_{0}^{g})$ grouped by goal category.
{tblr}
TEDUO: steps 1 & 2 + BabyAI-IL-bot train train 46 (± 2.0) 92 (± 1.2) 100 (± 0.0) 7 (± 2.9)
test train 30 (± 1.6) 58 (± 1.7) 100 (± 0.0) 6 (± 2.1)
train test 5 (± 1.2) 44 (± 2.5) 4 (± 0.9) 0 (± 0.0)
test test 7 (± 1.2) 40 (± 2.2) 5 (± 0.9) 0 (± 0.0)
TEDUO (Llama-3-8B) train train 46 (± 2.3) 85 (± 1.3) 100 (± 0.0) 0 (± 0.0)
test train 39 (± 2.4) 65 (± 2.0) 100 (± 0.0) 0 (± 0.0)
train test 20 (± 3.8) 87 (± 3.2) 83 (± 1.8) 0 (± 0.0)
test test 26 (± 3.0) 70 (± 2.8) 61 (± 2.6) 0 (± 0.0)
B.7 Generalization to Grids of Varying Sizes
In this experiment, we fine-tune Llama-3-8B-Instruct using the TEDUO pipeline, leveraging training data collected from environments with grid sizes 1, 4, 6, and 9. Here, the grid size refers to the number of rooms it contains. Figure B.3 illustrates an example of a 5$\times$5 grid (size 25). We evaluate the resulting model on larger environments with grid sizes 16 (4$\times$4), 25 (5$\times$5), and 49 (7$\times$7), as reported in Table B.7.
Classical CNN-based RL agents such as BabyAI-IL-bot fail to generalize to larger grids due to the fixed size of their convolutional layers, which constrains them to a fixed grid size. In contrast, the vanilla LLM exhibits consistent performance across grid sizes, albeit at a lower level. The model fine-tuned with TEDUO achieves substantially improved success rates on grid sizes close to those seen during training. This demonstrates TEDUO’s ability to facilitate generalization to more complex environments. However, as the grid size increases beyond the training distribution, the performance gradually declines, ultimately approaching that of the vanilla LLM on 7$\times$7 grids.
Table B.7: Online evaluation of generalization to larger grid sizes. Results are averaged over 200 $(g,s_{0}^{g})$ pairs.
{tblr}
width = colspec = Q[320]Q[129]Q[83]Q[142]Q[156]Q[100], cell21 = r=3, cell61 = r=6, hline1,12 = -0.08em, hline2 = -0.05em, hline5-6 = -, Method & Environ- ment Goals Success Rate Episode Length Invalid Actions
Llama-3-8B (vanilla) 4x4 train/test 13 (±0.8) 905 (±6.2) 33 (±0.2)
5x5 train/test 16 (±0.8) 901 (±6.3) 35 (±0.2)
7x7 train/test 19 (±0.8) 874 (±6.8) 39 (±0.1)
TEDUO: steps 1 2 + BabyAI-IL-bot 4x4 / 5x5 / 7x7 train/test 0 (±0) 500 (±0) NA
TEDUO-Llama-3-8B 4x4 train 53 (±1.8) 597 (±16.1) 33 (±0.6)
4x4 test 36 (±1.6) 726 (±14.3) 38 (±0.4)
5x5 train 35 (±1.8) 739 (±16.3) 36 (±0.6)
5x5 test 31 (±1.6) 757 (±13.3) 36 (±0.4)
7x7 train 19 (±1.3) 867  (±10.2) 31 (±0.4)
7x7 test 25 (±1.3) 852  (±10.1) 35 (±0.4)
Figure B.3: Example of a BabyAI grid of size 25 (5x5).
Appendix C Experimental details
C.1 Choice of Environments